# inkscapelatex=false

### v1

@@ 5. Auto-regressive self-attention with identity value matrix
for each $1\leq i\leq\ell$, $t\geq 0$ and $\mathbf{z}=({\rm z}_{1},\ldots,{\rm z}_{\ell})\in\mathbb{S}^{n}$, where $\beta_{ij}(t,\mathbf{z})=\alpha_{ij}(t,U^{-1}\mathbf{z})$, $1\leq i,j\leq\ell$. In other words, we obtain (15) with $h=1$ and $\beta_{ij}$ instead of $\alpha_{ij}$, which also satisfy Lemma 4.1. Therefore, Theorem 5.1 ensures that the consensus set $\mathcal{C}_{\ell}$ is asymptotically stable and the domain of attraction contains the set $\mathcal{D}_{\ell}^{1}$.
In the following section, we will investigate the stability of the system with non-identity value matrices while keeping the standard projection (3).
Remark 5.2 (Closest result available in the literature).
6. Symmetric value matrix
In this section we extend the results of the previous section to more general value matrices. To that end, we restrict ourselves to a single head and symmetric, time-independent value matrix, i.e., $h=1$ and $U_{1}(t)=U$ with $U^{\top}=U$. In addition, the tokens evolve on the sphere, i.e., $W=\mathbb{I}_{n+1}$, leading to:
(21) $\dot{\rm y}_{i}=T_{{\rm y}_{i}}\pi\cdot\left(\sum_{j=1}^{i}\alpha_{ij}(t,\mathbf{y})\,U\,{\rm y}_{j}\right)=\sum_{j=1}^{i}\alpha_{ij}(t,\mathbf{y})\left(U\,{\rm y}_{j}-{\rm y}_{i}^{\top}\,U\,{\rm y}_{j}\,{\rm y}_{i}\right),$
for each $1\leq i\leq\ell$, $t\geq 0$ and $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathbb{S}^{n})^{\ell}$, where:
$Z_{i}(t,\mathbf{y})=\sqrt{n+1}\,\sum_{j=1}^{i}\exp({\rm y}_{i}^{\top}\,P(t)\,{\rm y}_{j}),\quad\alpha_{ij}(t,\mathbf{y})=\left\{\begin{array}[]{ll}\displaystyle\frac{1}{Z_{i}(t,\mathbf{y})}\,\exp({\rm y}_{i}^{\top}\,P(t)\,{\rm y}_{j}),&i\geq j,\\ 0,&i<j.\end{array}\right.$
We denote the spectrum of $U$ by $\lambda(U)$. Note that $\lambda(U)\subset\mathbb{R}$ as $U$ is symmetric. Given $\lambda\in\lambda(U)$, the corresponding eigenspace is denoted by $L_{\lambda}(U)\subset\mathbb{R}^{n+1}$. Lastly, $L_{\lambda}(U)^{\perp}=\{w\in\mathbb{R}^{n+1}\mid w^{\top}\,v=0,~{}\forall v\in L_{\lambda}(U)\}$ denotes the orthogonal complement of $L_{\lambda}(U)$ (with respect to the Euclidean metric). Recall that $L_{\mu}(U)\subset L_{\lambda}^{\perp}(U)$ for each $\mu\in\lambda(U)-\{\lambda\}$.
Unlike the case $U=\mathbb{I}_{n+1}$ considered in the previous section, the first token is no longer fixed. However, it can be shown that it converges to a fixed position provided the multiplicity of the largest eigenvalue is one.
@@ 6. Symmetric value matrix
Let $\tilde{\alpha}:[0,\infty[\to[c,\infty)$ be a continuously differentiable function for some $c\in\mathbb{R}^{+}$ and let $\lambda=\max\lambda(U)$. Suppose that $\dim L_{\lambda}(U)=1$ and define $v\in\mathbb{S}^{n}$ by $L_{\lambda}(U)\cap\mathbb{S}^{n}=\{-v,v\}$. The only equilibria ${\rm y}^{*}\in\mathbb{S}^{n}$ of the following differential equation:
(22) $\dot{\rm y}=\tilde{\alpha}(t)\,(U\,{\rm y}-{\rm y}^{\top}\,U\,{\rm y}\,{\rm y}),\qquad{\rm y}\in\mathbb{S}^{n},~{}t\geq 0,$
@@
    ${\rm y}^{*}\in L_{\lambda}(U)\cap\mathbb{S}^{n}=\{-v,v\}$ is asymptotically stable with domain of attraction:
@@
Firstly, $L_{\mu}(U)\cap\mathbb{S}^{n}$ is a manifold of equilibria of (22) for each $\mu\in\lambda(U)$, since:
$\tilde{\alpha}(t)\,(U\,{\rm y}-{\rm y}^{\top}\,U\,{\rm y}\,{\rm y})=\mu\,\tilde{\alpha}(t)\,({\rm y}-{\rm y}^{\top}\,{\rm y}\,{\rm y})=0,\qquad{\rm y}\in L_{\mu}(U)\cap\mathbb{S}^{n},~{}t\geq 0.$
In order to study their stability, let ${\rm y}:[0,\infty[\to\mathbb{S}^{n}$ be a solution of (22) and define:
$b:[0,\infty[\to[-1,1],\quad t\mapsto b(t)=v^{\top}\,{\rm y}(t).$
From (22) and the symmetry of $U$, the dynamics of $b$ is readily seen to be:
@@
By using that ${\rm y}^{\top}\,U\,{\rm y}\leq\lambda\,{\rm y}^{\top}\,{\rm y}=\lambda$ (and the equality holds if and only if ${\rm y}\in L_{\lambda}(U)$), the fact that $\tilde{\alpha}(t)>0$ for each $t\geq 0$, we obtain the equilibria of the previous equation:
@@
Moreover, $\dot{b}<0$ for $b\in(-1,0)$ and $\dot{b}>0$ for $b\in(0,1)$. Hence, $b^{*}=0$ is unstable with empty domain of attraction, whereas $b^{*}=-1$ and $b^{*}=1$ is asymptotically stable with domains of attraction $[-1,0)$ and $(0,1]$, respectively, which corresponds $\mathcal{D}^{1}(-v)$ and $\mathcal{D}^{1}(v)$, respectively. ∎
The previous lemma settles the asymptotic stability of the first token and is used to prove the following result establishing asymptotic stability of two specific consensus points induced by the matrix $U$.
@@
Suppose that $U$ is symmetric, that $\dim L_{\lambda}(U)=1$, where $\lambda=\max\lambda(U)$, and denote the elements of $L_{\lambda}(U)\cap\mathbb{S}^{n}$ by $L_{\lambda}(U)\cap\mathbb{S}^{n}=\{-v,v\}$. If $\lambda>0$ and $\sup_{t\geq 0}\|P(t)\|<b$ for some $b>0$, then $\mathbf{y}^{*}=(v,\ldots,v)$ (resp. $\mathbf{y}^{*}=-(v,\ldots,v)$) is an asymptotically stable equilibrium of (21) and its domain of attraction contains the set:
$\mathcal{D}^{\ell}(v)=\{({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathbb{S}^{n})^{\ell}\mid v^{\top}\,{\rm y}_{i}>0,~{}1\leq i\leq\ell\},\qquad\text{(resp. $\mathcal{D}^{\ell}(-v)$)}.$
@@
Firstly, the dynamics of the subsystem of (21) given by the first token is given by $\dot{\rm y}_{1}=\tilde{\alpha}_{11}(t)(U\,{\rm y}_{1}-{\rm y}_{1}^{\top}\,U\,{\rm y}_{1}\,{\rm y}_{1})$, with $\tilde{\alpha}_{ij}:[0,\infty[\to[0,\infty[$ as in (14). Lemmas 4.1 and 6.1 ensure that the result holds for that subsystem.
Now we proceed by induction: given $2\leq i\leq\ell$, the result is assumed to hold for the subsystem of (21) given by the first $i-1$ tokens. Hence, for each solution $({\rm y}_{1},\ldots,{\rm y}_{i}):[0,\infty[\to(\mathbb{S}^{n})^{i}$ of the subsystem of (21) given by the first $i$ tokens, we have that $\lim_{t\to\infty}{\rm y}_{j}(t)=v$ for each $11\leq j\leq i-$ 1 provided $({\rm y}_{1}(0),\ldots,{\rm y}_{i-1}(0))\in\mathcal{D}^{i-1}(v)$. In order to study the behavior of the $i$-th token, we define:
$e_{j}:[0,\infty[\to\mathbb{R}^{n+1},$ $t\mapsto e_{j}(t)={\rm y}_{j}(t)-v,\qquad 1\leq j\leq i-1,$
$b_{i}:[0,\infty[\to[-1,1],$ $t\mapsto b_{i}(t)=v^{\top}\,{\rm y}_{i}(t).$
A straightforward check, using that $U$ is symmetric and $v\in\mathbb{S}^{n}$ as well as the previous definitions, leads to:
@@
(24) $=C_{2}\sum_{j=1}^{i-1}|e_{j}|,\qquad t\geq 0,$
with $c_{2}>0$ as in Lemma 4.1 and $C_{2}=c_{2}\,(\lambda+\|U\|)$, where we have used that $\lambda>0$.
Given that $\lambda>0$ and ${\rm y}_{i}^{\top}\,U\,{\rm y}_{i}\leq\lambda\,{\rm y}^{\top}\,{\rm y}=\lambda$, the only equilibria of (23) are $b_{i}^{*}=1$ (which corresponds to ${\rm y}_{i}^{*}=v$) and $b_{i}^{*}=-1$ (which corresponds to ${\rm y}_{i}^{*}=-v$). In order to analyze the stability of the former, let us consider the function $V_{+}:[-1,1]\to\mathbb{R}$ defined as $V_{+}(b_{i})=1-b_{i}$. We have that $V_{+}(1)=0$ and $V_{+}|_{(0,1)}>0$, as well as:
@@
where we have used (24), $\tilde{\alpha}_{ij}(t),\lambda>0$ for each $t\geq 0$, and $0<b_{i}\leq 1$. For $e=0$ and $t\geq 0$, $\dot{V}_{+}(t,1)=0$ and $\dot{V}_{+}(t,\cdot)|_{(0,1)}<0$. This ensures that $V_{+}$ is a strict ISS-Lyapunov function for the equilibrium $b_{i}^{*}=1$ of (23), where the input is given by $e=(e_{1},\ldots,e_{i-1})$. We conclude that $b_{i}^{*}=1$ is ISS-stable with domain of attraction $(0,1)$. This corresponds to ${\rm y}_{i}^{*}=v$ with domain of attraction $\{{\rm y}_{i}\in\mathbb{S}^{n}\mid v^{\top}\,{\rm y}_{i}>0\}$.
Lastly, if we regard the errors as function of the initial conditions i.e., $e(t)=e(t,{\rm y}_{1}^{0},\ldots,{\rm y}_{i-1}^{0})$, then the induction hypothesis ensures that $e^{*}=0$ is an asymptotically stable equilibrium and its domain of attraction contains $\mathcal{D}_{1}^{i-1}(v)$. As a result, $(e^{*},{\rm y}_{i}^{*})=(0,v)$ is asymptotically stable for the cascade system $(e,{\rm y}_{i})$ and its domain of attraction contains $\mathcal{D}_{1}^{i}(v)$ (cf. [33, Lemma 4.7]).
The proof for the case $\mathbf{y}^{*}=-(v,\ldots,v)$ is analogous. ∎

### v2

@@ 5.1. Identity value matrix
for each $1\leq i\leq\ell$, $t\in\mathbb{R}_{0}^{+}$ and $\mathbf{z}=({\rm z}_{1},\ldots,{\rm z}_{\ell})\in\mathbb{S}^{n}$, where $\beta_{ij}(t,\mathbf{z})=\alpha_{ij}(t,U^{-1}\mathbf{z})$, $1\leq i,j\leq\ell$. In other words, we obtain Eq.˜6 under ˜5.1 with $h=1$ and $\beta_{ij}$ instead of $\alpha_{ij}$, which also satisfy Lemma˜4.1. Therefore, Theorem˜5.1 ensures that the consensus set $\mathcal{C}_{\ell}$ is asymptotically stable and the domain of attraction contains the set $\mathcal{D}_{\ell}^{1}$.
5.2. Symmetric value matrix
Now we extend the results of the previous section to more general value matrices. As above, the tokens evolve on the sphere, i.e., $W=\mathbb{I}_{n+1}$.
Hypothesis 5.2.
The model Eq.˜6 is auto-regressive, $W=\mathbb{I}_{n+1}$, and we have:
  1. (1)
    There is only one head, i.e., $h=1$.
  2. (2)
    $U_{1}(t)=U$ with $U^{\top}=U$.
  3. (3)
    $P_{1}(t)=P(t)$ is bounded, i.e., $\sup_{t\in\mathbb{R}_{0}^{+}}\|P(t)\|<\infty$.
We denote the spectrum of $U$ by $\lambda(U)$. Note that $\lambda(U)\subset\mathbb{R}$ as $U$ is symmetric. Given $\lambda\in\lambda(U)$, the corresponding eigenspace is denoted by $L_{\lambda}(U)\subset\mathbb{R}^{n+1}$. Let us denote by $L_{\lambda}(U)^{\perp}=\{w\in\mathbb{R}^{n+1}\mid w^{\top}v=0,~\forall v\in L_{\lambda}(U)\}$ the orthogonal complement of $L_{\lambda}(U)$ (with respect to the Euclidean metric). Recall that $L_{\mu}(U)\subset L_{\lambda}^{\perp}(U)$ for each $\mu\in\lambda(U)-\{\lambda\}$.
Unlike the case $U=\mathbb{I}_{n+1}$ considered in the previous section, the first token is no longer fixed. However, the following result shows that it converges to a fixed position provided the multiplicity of the largest eigenvalue is one.
@@ 5.2. Symmetric value matrix
Let $\tilde{\alpha}:\mathbb{R}_{0}^{+}\to[c,\infty[$ be a continuously differentiable function for some $c\in\mathbb{R}^{+}$, and $\lambda=\max\lambda(U)$. If $\dim L_{\lambda}(U)=1$, then the only equilibria ${\rm y}^{*}\in\mathbb{S}^{n}$ of the following differential equation:
(17) $\dot{\rm y}=\tilde{\alpha}(t)(U{\rm y}-{\rm y}^{\top}U{\rm y}{\rm y}),\qquad{\rm y}\in\mathbb{S}^{n},~t\in\mathbb{R}_{0}^{+},$
@@
    ${\rm y}^{*}\in L_{\lambda}(U)\cap\mathbb{S}^{n}$ is asymptotically stable with domain of attraction:
@@
Firstly, $L_{\mu}(U)\cap\mathbb{S}^{n}$ is a manifold of equilibria of Eq.˜17 for each $\mu\in\lambda(U)$, since:
$\tilde{\alpha}(t)(U{\rm y}-{\rm y}^{\top}U{\rm y}{\rm y})=\mu\tilde{\alpha}(t)({\rm y}-{\rm y}^{\top}{\rm y}{\rm y})=0,\qquad{\rm y}\in L_{\mu}(U)\cap\mathbb{S}^{n},~t\in\mathbb{R}_{0}^{+}.$
In order to study their stability, let ${\rm y}:\mathbb{R}_{0}^{+}\to\mathbb{S}^{n}$ be a solution of Eq.˜17 and define:
$b:\mathbb{R}_{0}^{+}\to[-1,1],\quad t\mapsto b(t)=v^{\top}{\rm y}(t),$
where $v\in L_{\lambda}(U)\cap\mathbb{S}^{n}$. From Eq.˜17 and the symmetry of $U$, the dynamics of $b$ is readily seen to be:
@@
By using that ${\rm y}^{\top}U{\rm y}\leq\lambda{\rm y}^{\top}{\rm y}=\lambda$ (and the equality holds if and only if ${\rm y}\in L_{\lambda}(U)$), the fact that $\tilde{\alpha}(t)>0$ for each $t\in\mathbb{R}_{0}^{+}$, we obtain the equilibria of the previous equation:
@@
Moreover, $\dot{b}<0$ for $b\in]-1,0[$ and $\dot{b}>0$ for $b\in]0,1[$. Hence, $b^{*}=0$ is unstable with empty domain of attraction, whereas $b^{*}=-1$ and $b^{*}=1$ are asymptotically stable with domains of attraction [1,0[[-1,0[ and ]0,1]]0,1], respectively, which correspond to $\mathcal{D}^{1}(-v)$ and $\mathcal{D}^{1}(v)$, respectively. ∎
The previous lemma allows for establishing the asymptotic stability of two specific consensus points induced by the matrix $U$ using the same technique as in Theorem˜5.1. Namely, the dynamics of each other token ${\rm y}_{i}$, $2\leq i\leq\ell$, is projected to the real line using its inner product with the corresponding asymptotically stable equilibrium of ${\rm y}_{1}$. By following an induction argument and treating the distance of the previous tokens to the equilibrium as an error (as it converges to zero within time), an input-to-state-stability-Lyapunov function for the projected dynamics of ${\rm y}_{i}$ can be found, yielding the result.
@@
Suppose that ˜5.2 holds and $\dim L_{\lambda}(U)=1$, where $\lambda=\max\lambda(U)$. Let $v\in L_{\lambda}(U)\cap\mathbb{S}^{n}$. If $\lambda>0$, then $\mathbf{y}^{*}=(v,\ldots,v)$ is an asymptotically stable equilibrium of Eq.˜6 and its domain of attraction contains the set:
$\mathcal{D}^{\ell}(v)=\{({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathbb{S}^{n})^{\ell}\mid v^{\top}{\rm y}_{i}>0,~1\leq i\leq\ell\}.$
@@
Firstly, the dynamics of the subsystem of Eq.˜6 under ˜5.2 given by the first token reads: $\dot{\rm y}_{1}=\tilde{\alpha}_{11}(t)(U{\rm y}_{1}-{\rm y}_{1}^{\top}U{\rm y}_{1}{\rm y}_{1})$, with $\tilde{\alpha}_{ij}:\mathbb{R}_{0}^{+}\to\mathbb{R}_{0}^{+}$ as in Eq.˜11. Lemmas˜4.1 and 5.2 ensure that the result holds for that subsystem.
Now we proceed by induction: given $2\leq i\leq\ell$, the result is assumed to hold for the subsystem of Eq.˜6 under ˜5.2 given by the first i1i-1 tokens. Hence, for each solution $({\rm y}_{1},\ldots,{\rm y}_{i}):\mathbb{R}_{0}^{+}\to(\mathbb{S}^{n})^{i}$ of the subsystem of Eq.˜6 under ˜5.2 given by the first $i$ tokens, we have that $\lim_{t\to\infty}{\rm y}_{j}(t)=v$ for each $1\leq j\leq i-1$ provided $({\rm y}_{1}(0),\ldots,{\rm y}_{i-1}(0))\in\mathcal{D}^{i-1}(v)$. In order to study the behavior of the $i$-th token, we define:
$e_{j}:\mathbb{R}_{0}^{+}\to\mathbb{R}^{n+1},\qquad$ $t\mapsto e_{j}(t)={\rm y}_{j}(t)-v,\qquad 1\leq j\leq i-1,$
$b_{i}:\mathbb{R}_{0}^{+}\to[-1,1],\qquad$ $t\mapsto b_{i}(t)=v^{\top}{\rm y}_{i}(t).$
A straightforward check using that $U$ is symmetric and $v\in\mathbb{S}^{n}$, as well as the previous definitions, leads to:
@@
(19) $=C_{2}\sum_{j=1}^{i-1}|e_{j}|,\qquad t\in\mathbb{R}_{0}^{+},$
with $c_{2}>0$ as in Lemma˜4.1 and $C_{2}=c_{2}(\lambda+\|U\|)$, where we have used that $\lambda>0$.
Given that $\lambda>0$ and ${\rm y}_{i}^{\top}U{\rm y}_{i}\leq\lambda{\rm y}^{\top}{\rm y}=\lambda$, the only equilibria of Eq.˜18 when $e=0$ are $b_{i}^{*}=1$ (which corresponds to ${\rm y}_{i}^{*}=v$) and $b_{i}^{*}=-1$ (which corresponds to ${\rm y}_{i}^{*}=-v$). In order to analyze the stability of the former, let us consider the function $V_{+}:[-1,1]\to\mathbb{R}$ defined as $V_{+}(b_{i})=1-b_{i}$. We have that $V_{+}(1)=0$ and $V_{+}|_{]0,1[}>0$, as well as:
@@
where we have used Eq.˜19, $\tilde{\alpha}_{ij}(t),\lambda>0$ for each $t\in\mathbb{R}_{0}^{+}$, and $0<b_{i}\leq 1$. For $e=0$ and $t\in\mathbb{R}_{0}^{+}$, $\dot{V}_{+}(t,1)=0$ and $\dot{V}_{+}(t,\cdot)|_{]0,1[}<0$. This ensures that $V_{+}$ is a strict ISS-Lyapunov function for the equilibrium $b_{i}^{*}=1$ of Eq.˜18, where the input is given by $e=(e_{1},\ldots,e_{i-1})$. We conclude that $b_{i}^{*}=1$ is ISS-stable with domain of attraction ]0,1]]0,1]. This corresponds to ${\rm y}_{i}^{*}=v$ with domain of attraction $\{{\rm y}_{i}\in\mathbb{S}^{n}\mid v^{\top}{\rm y}_{i}>0\}$.
Lastly, if we regard the errors as function of the initial conditions i.e., $e(t)=e(t,{\rm y}_{1}^{0},\ldots,{\rm y}_{i-1}^{0})$, then the induction hypothesis ensures that $e^{*}=0$ is an asymptotically stable equilibrium and its domain of attraction contains $\mathcal{D}_{1}^{i-1}(v)$. As a result, $(e^{*},{\rm y}_{i}^{*})=(0,v)$ is asymptotically stable for the cascade system $(e,{\rm y}_{i})$ and its domain of attraction contains $\mathcal{D}_{1}^{i}(v)$ (cf. Lemma 4.7 of [34]). ∎